The aircraft, whose engine power is 3 MW, under the action of the traction force 4.5 kN flies 600 km for 0.5 hours. Determine the efficiency of aircraft engines.


1
Expert's answer
2021-04-19T17:08:43-0400

Find the speed:


"v=\frac dt."

Find power input:


"P_i=Fv=\frac{Fd}{t}."

Power output:


"P_o=3·10^6\text{ MW}."

Efficiency:


"\eta=\frac{P_i}{P_o}=\frac{Fd}{P_ot}=0.5"